"a stone is throwen in a vertically up wards directionwith the velocity of 5 m/s.if the acceleration of the stone is 10m/s in the downwords direction .what will be the height attained by the stone and how much time will it take t
 

Grade:9

1 Answers

Gunjan Rawal
9 years ago
initial velocity= 5 m/s
acceleration acts downwards that is the gravitational acc.  therefore the acc. of the body is  (-g)
so, now when the stone will reach the highest point it’s final velocity will be zero  .. (v = 0 ) 
now applying kinematics formulae  v2-u2=2.g.s   (s=displacement or max. height )
-(5)2=-2.10.s     therefore s = 5/4 m or 1.25 m .. now for time taken use formulae v = u + at 
therefore 0 = 5 – 10t  ==> 5=10t  ==> t = 0.5s ..!! answer. Height = 1.25m and Time taken = 0.5s